OpenStudy (anonymous):

The area of a rectangular garden is given by the trinomial x2 + 6x – 27. What are the possible dimensions of the rectangle? Use factoring. A. x – 9 and x + 3 B. x + 9 and x – 3 C. x – 9 and x – 3 D. x + 9 and x + 3

OpenStudy (zehanz):

If you factor, you get (x+a)(x+b), where a+b=6 (coefficient of x), and ab=-27 (constant term. Try it now..

OpenStudy (anonymous):

i think its A?

OpenStudy (zehanz):

-9+3=-6

OpenStudy (anonymous):

so its not a?

OpenStudy (zehanz):

Nope. Try again? a+b=6 ab=-27

OpenStudy (anonymous):

x + 9 and x + 3?

OpenStudy (zehanz):

9+3=12 instead of -6, 9*3=27 instead of 27

OpenStudy (zehanz):

I would take -3 and 9, so (x-3)(x+9), which would be B. Check: -3+9=6 OK -3*9=-27 OK See? That's all!

OpenStudy (anonymous):

i am confusedd

OpenStudy (anonymous):

r u sayying its x + 9 and x – 3

OpenStudy (zehanz):

You don't have to be, because it is entirely clear: you are looking for two numbers, a and b, to write your trinomial as (x+a)(x+b) If you would expand (x+a)(x+b), you would get (after tidying up): x²+(a+b)x+ab. Now you have: x²+ 6x -27 That is why I say: a+b=6 and a*b=-27 If you think a while about this, you'll come up with a=-3 and b=9.

OpenStudy (zehanz):

So, because x²+6x-27=(x-3)(x+9) and area=length*width, the dimensions can only be x-3 and x+9

OpenStudy (anonymous):

okay!!! haha i get itt. thank you.

OpenStudy (zehanz):

YW!